?
Competitive Exams MCQs
Aptitude
English
Verbal Reasoning
Non Verbal Reasoning
General Knowledge(GK)
General Science(GS)
Data Interpretation(DI)
Banking Awareness
Current Affairs
Computer Fundatmental MCQs
Computer Fundamentals
Microsoft Word
Microsoft Excel
MS Power Point
Tally
Operating System
Python
AI(Artificial Intelligence)
Join the Discussion
Login to post a comment or share your explanation.
Loginnow no detection mechanism is present
so per loss = total_bits_transmitted *probability(error)
loss = 9*0.001=0.009
Probability that single bit is in error = 10^-3 = 0.001
Single bit not in error = 1 - 0.001 = 0.999
Chances that all bits are correct = 0.999^9 = 0.991
Residual error rate = Channel that at least one of 9 bits is incorrect = 1 - 0.991 = 0.009